Minutes of the 2nd Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson Minutes taker Mark Dennehy Present Steven Watterson (Captain),

Benjamin Jones (Treasurer),

Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer),

Absent(with apologies) Conor McDermottroe (Public Relations Officer),

Absent Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 18 June 2001 1700 – 2000 Distribution DURC Committee

Overview of topics

1 Reading of Minutes

1 Reading of Minutes 1 1.1 Tasks from last meeting . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Removal Orders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Fassaroe Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Summer Shooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.5 Supervision of Shooters . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.6 Range improvements . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.7 50m Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.8 Bank Fees . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.9 rifle.club@tcd.ie . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.10 Coaching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.11 BUSA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.12 3P Air Rifle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.13 DUCAC UK funding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.14 ISS Bill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.15 Social Aspect of club . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.16 Inventory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.17 Ammunition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2 Johns Rifle 6

3 Secretarys Report 6

4 Treasurers Report 6

5 CRO report 6

6 A&E Report 6

7 TTT 7

8 Comber Squad Training weekend 7

9 50m Training 7

10 Any Other Business 7 10.1 Buddy System Trial . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 10.2 Bulldog Clips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 10.3 Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 10.4 Next Meeting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 10.5 Social event . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

1 Reading of Minutes

The minutes were taken as read, though Mark pointed out that two content changes had been made, as he had forgotten to note that Ben had handled the changing of signatories on the DURC account and he had also mistaken the notes regarding inquiries into DUCAC funding of UK trips.

Matters arising took some time however...

1.1 Tasks from last meeting

Little progress made here, things being hampered by lack of knowledge of who the new UCDRC committee is and Geoff Cooney and Pat Moray being on holidays.

1.2 Removal Orders

Alan suggested that we get removal orders in batches - Mark pointed out that the calendar was in chaos at present. Alan to get removal orders for the Comber squad training weekend and the 50m Nationals. Task: Removal orders for Comber squad training and 50m Nationals (Alan)

1.3 Fassaroe Open

Steven gave Ben receipts for the Fassaroe Open fees and ammunition bought there. Ben expressed regret about the purchase of Super Club ammunition, as not many people like it. Ben collected cash from Mark, John, Alan. Paying for competitions in future was discussed - it was agreed that we would ask for an invoice to be sent to DURC, which would then be paid promptly; but that Range Officers would collect cash from the DURC shooters before we left for the competition.

1.4 Summer Shooting

There was a spirited discussion on the topic of summer shooting arrangements. John was not happy at the thought of formal arrangements for ROing for people during the summer - he stated that the informality of shooting during the summer was the ROs reward for the termtime work they do. Various solutions were discussed with different levels of formality. Workload on the RO who was shooting that night was a problem identified, along with the idea of one person doing paperwork for several ROs as well as non-ROs, which was felt to be unfair. The convienence of giving notice to people was another problem. Steven volunteered to coordinate the arrangements for shooting and (after Alan wondered if that was the CROs job) that was accepted. John asked for it to be noted that informality was the key word in these arrangements. Task: Coordinate summer shooting arrangements (Steven)

1.5 Supervision of Shooters

This was discussed in light of the events at the Fassaroe Open where Joe Lalor had accidentally shot the wrong course of fire during the scottish match. John said this illustrated the need for a “nanny” at competitions even with small numbers of competitors - there had been three present at the time. Mark mentioned the “buddy” system from the previous meeting and this was discussed, John pointing out that it could not have worked at the FSC open as all three shooters had to shoot together. Steven agreed that in such cases it would not work, but that it worked in normal circumstances in Edinburgh. Mark pointed out that it was highly unusual for all shooters to start together, and that we did not have the equipment for that to happen normally, nor did ranges have enough firing points for it. It was agreed that the buddy system was a good idea but that a single person had to act as team manager, ie. booking details, buying ammunition, arranging payment of fees, etc. It was later agreed that the upcoming 50m National Championships would serve as a trial for this system and that it would be properly organised this time as the FSC open had not been properly setup this way.

Alan proposed a fast explaination be given by the nanny to the shooters of the course of fire for the competition before leaving the range that morning; John suggested that one target of each type be put up on the wall in the club to facilitate this.

Alan proposed a set of guidelines for Range Officers at competitions; he agreed to prepare these.

It was agreed that more information was needed on who was going to show up at the competitions; John suggested not broadcasting our departure time from the range except to those who said they would be attending. Task: Prepare set of guidelines for ROs at competitions (Alan)

1.6 Range improvements

Alan is going to put in a Halogen floodlamp at the 25yd targets to try to improve the lighting.

John suggested modification of the backstop to allow the two target cards to be mounted closer to the center of the range. This was to improve the sight picture on the targets, which is currently encroached on by the side walls. Mark said that the easiest route was to totally replace the current backstop with a new one, which would require cutting out some brickwork. This seemed practicable, but locating an 8’×3’ sheet of mild steel seemed difficult. Asking the Buildings Office was considered, but the odds of them doing it seemed long.

1.7 50m Open

John reported that the MRC wanted £5 per head on the day, and would only let us shoot on 10 of the 12 points, the other two being used for fullbore. MRC had said they had installed soundproofing, but skepticism was expressed on the effectiveness of this. Mark suggested letting MRC know that this was totally unacceptable before rejecting the idea totally. The donation

was actually £200.

FSC was then discussed as a venue - John was adamant that no range hire be paid as we had paid them £300 last year as a donation. This donation was discussed. Steven wondered how many years of goodwill it had bought, John estimated two, but would have preferred twenty, and suggested we milk it. Task: Give MRC a last chance to accept our terms before dropping them as a venue for the Open (John)

1.8 Bank Fees

Ben reported that AIB now wanted to charge us bank fees for writing sterling cheques and that he had pointed out that we were a student club, and that if they persisted, DURC would use ulster bank for foreign drafts.

1.9 rifle.club@tcd.ie

Mark reported no progress on getting this redirected to the committee list - John pointed out that it gets a lot of spam email. Steven enquired about filtering, but no adaquate solution was found.

1.10 Coaching

Mark reported Rhona was okay with the idea of doing a nights coaching. He suggested that we try to get more money to properly pay for coaching, Ben said that DUCAC will not pay for coaches, only cover their expenses. Mark said he would ask more national squad members as the opportunity arose. Task: Continue to inquire into Squad members coaching (Mark)

1.11 BUSA

Steven agreed to follow up on the BUSA entries as Marks to-do list was looking long. Task: Follow-up BUSA entries (Steven)

1.12 3P Air Rifle

A lenghty and confused discussion was held on the 3P Air rifle proposal. Mark reported that Ray Kane of WTSC thought it was an excellent idea and that WTSC might be interested in trying 3P that way. Ben and Mark then noted that that meant we would have to start training. The discussion of the firing points arrangements then got very confusing, with everyone having different mental pictures of what was being discussed. Eventually, it was agreed that the best way to implement the idea was to have a target frame constructed for each firing point. This frame, similar to FSC’s target frames, would be freestanding, approx. 6’×2’, and would have three points for mounting air targets on it, one for each position. This would be a simple construction job, ideally suited to Keith White’s DIY talents. Task: Design and build 3P Air Rifle target holders (Unassigned)

1.13 DUCAC UK funding

How to go about discreet enquiries was discussed: the best approach was felt to be asking around at the DUCAC procedures meeting in October.

1.14 ISS Bill

John asked if this was really needed and if we can afford it. Ben said we had no problem paying it this year and that it is normally taken from hire and that that category in the grant normally has only one other entry, that of the WTSC hall for the colours competition. Steven asked what the network link was used for: Mark said that it allowed full control of content on the website and more advanced things that we couldn’t do on alf2; that it allowed for a central point for email such as captain@durc.tcd.ie. John pointed out that this was served by rifle.club@tcd.ie. Mark said that the main reason he wanted the network link was to allow the RCMS system to work. Steven asked if online booking was useful or not, Mark pointed out that we had no data to gauge that on but that in a larger new range, we would be swamped with the current paperwork - Steven pointed out that the work on the new Sports Center appeared to be well behind schedule. It was decided not to pay the bill until asked by ISS; and that it should be continued for the next academic year.

1.15 Social Aspect of club

This merged with the previous topic. Playstation was discussed, with Steven and Mark unhappy at games being played on it, while Alan and John felt it kept people on the range after shooting. No decision was made on the situation. Free Tea and Coffee was accepted as an excellent idea, but there was a minor debate on biscuits as their consumption was very high.

The £600 spent on beer was thoroughly questioned by John, who felt we had spent less this year on beer than any other year. Ben spent some time tracing the amounts in the accounts. Mark pointed out that while the amount of cash seemed far too high to spend on beer anyway, that the main problem had been the unscheduled nature of beer purchases. Steven said that the obvious solution was to budget for beer purchases instead of not allowing for this expense. Ben pointed out that the clubs low income did make this difficult, but all agreed it was the correct solution. Task: Budget for Beer (Unassigned)

1.16 Inventory

John asked Mark for a list of items he had bought the club at Munich. Task: Email John a list of club items bought at Munich (Mark)

1.17 Ammunition

John reported that he had been looking at this problem as A&E Officer and had made progress. Dave Cooney’s supplier had agreed to sell ammuntion to us directly with savings of 12-15%. The main problem was that Dave would hear of this and might react badly. This was felt not to be sufficent reason not to pursue this directly. Quantities would be a problem as the supplier will only deal in levels of 5-10,000 rounds per order. Mark suggested raising the maximum amount of ammunition the club had on its licence to allow for us to buy these quantities in several grades of ammunition; John said that the limit was Pat Moray saying that we had only so much room for storage and agreed to take that up with him at a later date. John reported that Federal, Eley and Lapua had not responded to his correspondance. Steven suggested contacting RWS as well. Steven also suggested asking if we could get Pellets for air rifle from the new supplier as if Dave decided not to act as our gunsmith as a result of losing our business we would need a new supply. John agreed to follow up and find if different grades of .22 and Air pellets would be available.

2 Johns Rifle

Mark asked about ammunition profiles, ie. the grades of ammunition we bought. John said that if the new accounting system was adopted by Pat Moray then it would be no problem to do so and we should discuss it at that point as batch testing was a difficult task.

John reported that he had no problems with Pat Moray and as Mark agreed, that it must just be that Pat and Jason did not get along well. Task: Follow up on new ammunition supplier (John)

2 Johns Rifle

It was decided to number Johns rifle as #9, as it did not have any number on it at present.

3 Secretarys Report

Mark ran through the correspondance sent since the AGM. It was decided not to use registered post for postal cards due to expense and instead to scan or photocopy them and get proof of postage. Mark replied that he had registered the NSRA cards as they were late and that since only Ben had shot his on time that this was a necessary measure at the time.

4 Treasurers Report

Less than £200 remain in our account. Ben has prepared the rules regarding financial dealings with DURC and Mark is to give these to Sandy Wilson at the next NTSA committee meeting. The NSRA invoice has been paid (for the range plans). Task: Give financial rules to Sandy Wilson (Mark)

5 CRO report

Alan is continuing to tidy the range through delegation but will hold off difficult tasks till september so that the range won’t be cleaned and then dirtied before term started. Alan will put in a Halogen floodlamp and get removal orders for Comber and the 50m Nationals. Task: Tidy range (Alan) Task: Install halogen floodlight on range (Alan) Task: Get removal orders for Comber and 50m Nationals (Alan)

6 A&E Report

Mark asked that the A&E officer report on the ammunition levels for each committee meeting; when John complained that this was akward Mark reminded him that he’d

7 TTT

asked for it! Mark also noted we needed at least 1000 rounds of Lapua Master for the

Comber weekend.  1905,  John

leaves to catch a bus

7 TTT

Ben asked if the deadline for the entries was this friday and how many had we: Mark said it was this friday, we had one in from the Dublin Rifle Club and that RRPC had been in touch about entries. Mark said he was going to make some reminder phone calls this week to the various clubs.

Steven asked what clubs Mark had invited in Scotland and the Isle of Man; but said that it was unlikely that Derek Wilde had passed on entry forms and so forth as he was a pistol shooter.

Steven wondered if the cards arriving each fortnight would annoy DUCAC; Mark said he’d mention it to Cathy but that he thought it would not be a problem. Task: Contact more UK clubs about the TTT (Steven) John returns,

his bus being later than thought

8 Comber Squad Training weekend

The question of who was driving was a problem as Mark’s licence is void in N.Ireland and his car is not on open insurance. Mark suggested getting a lift from members of the Squad who were going or putting gear in Matt Foxs van if WTSC were taking it. As we weren’t sure if he was going, it was decided that Mark would contact Comber to find out about transport and to follow up on arrangements. Task: Arrange transport to/from Comber range (Mark)

9 50m Training

John leaves

Days for training in FSC for 50m were discussed. Mark suggested he would attempt _{again.} to get a parking permit for TCD, bring in his car in the morning and then depart for FSC at 1500 on fridays. This seemed a good idea, but Mark pointed out that as Dave Cooney was on the FSC committee and we were about to drop him as Ammunition supplier that we might have to pay an affiliation fee. John didn’t like the idea as we’ve had an agreement with FSC that DURC members could train on their range for some time. Agreed to go training in FSC on the weekend before the Nationals : 29/06-1/07

10 Any Other Business

10.1 Buddy System Trial

Mark proposed the 50m Nationals serve as a full trial for the new buddy system for supervising shooters at competition. This was agreed.

10 Any Other Business

10.2 Bulldog Clips

Steven noted that we had insufficent clips for 50m training. The estimate was that we should have enough for five people to train with: this came to 18 clips as we had enough for two people allready. Mark suggested putting six clips on a string and thus keeping them together in single-firing-point units. Ben agreed that we could buy them but suggested we try other means first. Steven agreed to follow up. Task: Acquire bulldog clips (Steven)

10.3 Tools

Mark slapped Alan for removing the walther’s foresight hex key from the kk200 sights box. Alan apologised and was worried that it hadn’t been in the other toolboxes. Mark pointed out that this highlighted the need for a single toolbox that could be taken to competitions as otherwise tools would be lost or misplaced which could ruin a competition - Mark nearly having to forego shooting at the FSC Open as a result of this.

10.4 Next Meeting

The scheduling of the next meeting was discussed. Ben doesn’t like the Common Room as it’s open, noisy and other people use it during meetings. Mark said he’d found no rooms where we wouldn’t be charged for a security guard to lock up. Alan said that he had access to the seminar room in the Parsons Building and it was agreed to try there for the next meeting. As Alans results were due out the following monday, the next meeting was scheduled for 1800h on Wednesday June 27th. The agenda is to include the DURC Open and the Calendar for the coming years events.

10.5 Social event

It was noted that some of the committee were going to shoot and then go for a drink afterwards on Thursday evening, invitations were extended to all present.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

Tasks agreed on at meeting

Alan Removal orders for Comber squad training and 50m Nationals (18

June 2001) [/149] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Steven Coordinate summer shooting arrangements (18 June 2001) [/179] . . /2 Alan Prepare set of guidelines for ROs at competitions (18 June 2001) [/217] /2 John Give MRC a last chance to accept our terms before dropping them

as a venue for the Open (18 June 2001) [/256] . . . . . . . . . . . . . . /2

Mark Continue to inquire into Squad members coaching (18 June 2001) [/280] /2 Steven Follow-up BUSA entries (18 June 2001) [/287] . . . . . . . . . . . . /2 Unassigned Design and build 3P Air Rifle target holders (18 June 2001) [/306] . /2 Unassigned Budget for Beer (18 June 2001) [/356] . . . . . . . . . . . . . . . . . /2 Mark Email John a list of club items bought at Munich (18 June 2001) [/363] /2 John Follow up on new ammunition supplier (18 June 2001) [/398] …. /2 Mark Give financial rules to Sandy Wilson (18 June 2001) [/419] . . . . . /2 Alan Tidy range (18 June 2001) [/430] . . . . . . . . . . . . . . . . . . . . /2 Alan Install halogen floodlight on range (18 June 2001) [/431] . . . . . . . /2 Alan Get removal orders for Comber and 50m Nationals (18 June 2001) [/432] /2 Steven Contact more UK clubs about the TTT (18 June 2001) [/461] … /2 Mark Arrange transport to/from Comber range (18 June 2001) [/476] . . /2 Steven Acquire bulldog clips (18 June 2001) [/512] . . . . . . . . . . . . . . /2